OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 4) with a slope of –1 in standard form.

OpenStudy (anonymous):

y=mx+b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

use the equation y - y1 = m(x - x1)

OpenStudy (anonymous):

y - 4 = -1(x - -1) this is the equation but you can put it in point slope form now

OpenStudy (anonymous):

here are the choices: a. –x + 4y = –1 b. x – 4y = –1 c. x + y = 3 d. x – y = –3

OpenStudy (anonymous):

Yeah I just read it wants it in standard form so y - 4 = -1(x - -1) y -4 = -1x - 1 y =-1x - 1 +4 y= -1x + 3 <---slope point form and now transform to standard form 1x + y = 3 answer = c

OpenStudy (anonymous):

thanks :)